OpenStudy (erinweeks):

2. (2.05) Given (3, -2) and (-2, 3), a) Write the equation of the line in slope-intercept form that passes through these two points. You will be asked to explain your answer. b) What is the slope of any line parallel to this line? Why? c) What is the slope of any line perpendicular to this line? Why?

OpenStudy (erinweeks):

@rajee_sam @jim_thompson5910

OpenStudy (phi):

Did you find the slope ?

OpenStudy (erinweeks):

what I got in the end was y=5/-5x-4

OpenStudy (erinweeks):

m=y2-y1/x2-x1 m=3-(-2)/-2-3 m=5/-5 y-y1=m(x-x1) y-(-2)=5/-5(x-3) y+2=-1(x-3) -2 -2 y=5/-5x-4 but I know this is wrong

jimthompson5910 (jim_thompson5910):

what does 5/(-5) reduce to

jimthompson5910 (jim_thompson5910):

whenever you're dealing with fractions, you should always reduce as much as possible.

OpenStudy (erinweeks):

Now what I have from this is m=y2-y1/x2-x1 m=3-(-2)/-2-3 m=5/-5 y-y1=m(x-x1) y-(-2)=5/-5(x-3) y+2=-1(x-3) -2 -2 y=-1x+1

jimthompson5910 (jim_thompson5910):

you got it, your equation in slope intercept form is y = -1x+1 or y = -x+1

OpenStudy (erinweeks):

But she said in the comments that that like had to be changed to this that I had y+2=-1(x-3) and then she stated that I needed to say the slopes

OpenStudy (erinweeks):

im confused on how to get the slope Jim all I have left is to state the slopes

jimthompson5910 (jim_thompson5910):

you corrected some of your work, but I still see 5/-5 in there

jimthompson5910 (jim_thompson5910):

when it should be -1

jimthompson5910 (jim_thompson5910):

on part b), what is the slope of any parallel line to this given line?

OpenStudy (rajee_sam):

no finally she wrote as -1x + 1

OpenStudy (erinweeks):

where would I put the -1 in there for this cause my teacher said only one line should be changed or something ? m=y2-y1/x2-x1 m=3-(-2)/-2-3 m=5/-5 y-y1=m(x-x1) y-(-2)=5/-5(x-3) y+2=-1(x-3) -2 -2 y=-1x+1

jimthompson5910 (jim_thompson5910):

this line y-(-2)=5/-5(x-3) should be this y-(-2)=-1(x-3)

OpenStudy (erinweeks):

and is the slope of this equation since its in slope intercept form -1x/1?

OpenStudy (rajee_sam):

you have everything right there is nothing to be changed here

jimthompson5910 (jim_thompson5910):

and m=5/-5 should be something like m=5/-5 = -1 or just m = -1

OpenStudy (rajee_sam):

The first part of the question wants her to write it in slope - intercept form which she has y = -1x + 1

OpenStudy (erinweeks):

ok is my equation correct now? and is the slope I stated correct or no? m=y2-y1/x2-x1 m=3-(-2)/-2-3 m=5/-5 y-y1=m(x-x1) y-(-2)=-1(x-3) y+2=-1(x-3) -2 -2 y=-1x+1

OpenStudy (erinweeks):

Slope is: -1x/1?

OpenStudy (rajee_sam):

slope is just the no. in front of x which is -1

OpenStudy (erinweeks):

so the slope would just be -1? not -1/1? Cause i thought slopes were usually in fraction form

jimthompson5910 (jim_thompson5910):

because you can write -1/1 as -1, I would write the slope as -1 yes that's true you usually see the slope as a fraction (eg: 2/3), but if you can go from a fraction to a whole number, then it makes things a whole lot easier

OpenStudy (erinweeks):

SO would the reciprocal slope be just positive 1?

OpenStudy (rajee_sam):

when you are graphing the line using slope and y-intercept for convenience sake you write it as fraction if it is not one. But all other purposes if the slope is nota fraction you write it as a whole number.

jimthompson5910 (jim_thompson5910):

so you cannot go from 2/3 to a whole number, so you leave it as 2/3 but you can go from -1/1 to -1, so I would stick with -1 and yes, the negative reciprocal of -1 is +1 or just 1

OpenStudy (erinweeks):

so the flipped one is just positive 1? Thanks Jim :) almost done I have one more section of this oral to complete then im done can you help me real fast cause I have to start gettting ready for work lol!

OpenStudy (phi):

remember a perpendicular line 's slope is the negative reciprocal that means multiply by -1 then flip: you start with m= -1 negate it : -(-1) =+1 flip it: 1/+1 which is just 1 so the negative reciprocal of -1 is +1